# An Efficient Learning Control Framework With Sim-to-Real for String-Type Artificial Muscle-Driven Robotic Systems

### v2

@@ V-A Training Results
To evaluate the effectiveness of the proposed bootstrap and augmentation techniques, we conducted an ablation study on two robotic systems. Training results of our learning control design, with and without bootstrap and augmentation, are presented in Fig. 6. The policies were trained and evaluated in simulation with random dynamics. As illustrated in the figures, our proposed algorithm exhibits the highest level of data efficiency in both robotic systems. Compared to the SAC algorithm trained solely with dynamics randomization (referred to as ”No BA”), the integration of bootstrap and augmentation enhances data efficiency and expedites the training process. In the robotic eye system, the baseline algorithm achieves an average reward of $-0.10$ after $2000$ training episodes, whereas our approach achieves comparable performance in only approximately $750$ episodes, resulting in a data efficiency increase to $2.67$ times the original level. Similarly, in the robotic wrist system, the baseline algorithm achieves an average reward of $-1.61$ after $3500$ episodes, while ours achieves comparable performance in only about $1200$ episodes, leading to a data efficiency increase to $2.92$ times. Notably, in the case of the more complex robotic wrist system, the demonstrations provided by the PID controller yield a more significant acceleration effect during the early stages of training. Meanwhile, the augmentation method demonstrates a favorable impact on enhancing the training speed in both systems.
(a) vanilla-SAC
(b) SAC-R
(c) SAC-BAR
(d) Box-and-Whisker Plot
Figure 9: Experimental results of the steady-state error tests in the parallel robotic wrist under selected control algorithms. (a), (b), (c) Heat maps illustrating the experimental results for the vanilla SAC algorithm, SAC with the randomization of muscles only (SAC-R), and our complete efficient learning control algorithm (SAC-BAR), respectively. The combination of target angles $(\theta^{*},\phi^{*})$ is systematically spaced at $2.5^{\circ}$ intervals within the range $(|\theta^{*}|,|\phi^{*}|)\leq 10^{\circ}$. The color bars in the visualization indicate the Euclidean distance norm of steady-state errors. (d) The box charts illustrating the medians, maxima, minima, quartiles, and outliers of the steady-state errors for the selected comparative control algorithms.
V-B Ablation Study in Robotic Wrist Experiments
To evaluate the impact of each component on real-world control performances, three different control policies are tested on the robotic wrist platform, including the vanilla SAC algorithm, SAC with only randomization of muscles (SAC-R), and our proposed efficient learning control framework (SAC-BAR). All three algorithms undergo training for an equivalent number of episodes in simulation, employing identical hyperparameters as specified in Table IV. During each experimental episode, the parallel robotic wrist starts from the initial state with $\theta=\phi=0^{\circ}$ and attempts to reach and maintain the target angle set $(\theta^{*},\phi^{*})$. The action vector $a$ is computed by the trained actors at a time step of $t_{a}=0.5$s, and each episode spans a duration of $t_{\text{total}}=20$s. Observations are measured by an IMU positioned at the center of the upper plate.
TABLE V: Averaged steady-state error over a total of 81 targets in ablation study with the robotic wrist
Agent Vanilla-SAC SAC-R SAC-BAR
Mean error (deg) $0.911\pm 0.457$ $0.657\pm 0.386$ $0.580\pm 0.368$
The experimental trajectories for two pairs of desired target angles, $(\theta^{*},\phi^{*})=(-5^{\circ},7.5^{\circ})$ and $(\theta^{*},\phi^{*})=(2.5^{\circ},-7.5^{\circ})$, of the vanilla SAC algorithm and our proposed SAC-BAR are illustrated in Fig. 7 and Fig. 8, respectively. In Fig. 7LABEL:sub@subfig:w_tra1_a and LABEL:sub@subfig:w_tra1_c, it is evident that without randomization of muscle dynamics, the vanilla-SAC agent struggles to achieve the desired orientation. While the agent may eventually reach the desired orientation at certain target angles, maintaining that target angle becomes a challenge, resulting in a considerable steady-state error. Conversely, as observed in Fig. 8LABEL:sub@subfig:w_tra2_a and LABEL:sub@subfig:w_tra2_c, the proposed SAC-BAR agent achieves the target state in less than $10$s and consistently remains within its neighborhood thereafter. This success is attributed to the agent learning to maximize the action value during the rising phase and promptly reduce it upon approaching the target angles, as illustrated by the input action trajectories in Fig. 8LABEL:sub@subfig:w_tra2_b and LABEL:sub@subfig:w_tra2_d. Additionally, when overshoot occurs, the agent still learns to return to and maintain the target orientation. Conversely, in Fig. 7LABEL:sub@subfig:w_tra1_b and LABEL:sub@subfig:w_tra1_d, there exist serious oscillations in the action trajectories of vanilla SAC agent during the initial $10$s, leading to a prolonged rise time and sometimes even failure to reach the target.
@@ V-B Ablation Study in Robotic Wrist Experiments
Figure 10: Experimental trajectories of the foveation test of the 2-DOF robotic eye. Target angles are set at $(\theta^{*}=-5^{\circ},\psi^{*}=7.5^{\circ})$ and $(\theta^{*}=10^{\circ},\psi^{*}=-5^{\circ})$, respectively. (a), (c) The system output angles $(\theta,\psi)$. (b), (d) The corresponding actions $a_{t}$ generated by the trained actor.
In order to comprehensively evaluate the control performance of the algorithms across different target angles, we design a steady-state error field test for the robotic wrist. We select a series of target angles $(\theta^{*},\phi^{*})$ spaced at $2.5^{\circ}$ intervals within the defined range $\theta^{*},\phi^{*}\in[-10^{\circ},10^{\circ}]$. Each agent is set to run for $25$s, sequentially targeting each angle pair. The steady-state error is measured and averaged over the last $5$s window, determined by computing the Euclidean distance norm between the current and target angles using $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\phi-\phi^{*})^{2}}$.
The steady-state errors of the field test of vanilla-SAC, SAC-R, and SAC-BAR are presented in Fig. 9LABEL:sub@subfig:a, LABEL:sub@subfig:b, and LABEL:sub@subfig:c, respectively. The $x$ and $y$ axes therein represent $\theta^{*}$ and $\phi^{*}$, while the color bars represent the averaged steady-state errors. Following that, box charts in Fig. 9LABEL:sub@subfig:d illustrate the medians, maxima, minima, quartiles, and outliers of the steady-state errors for different control policies. The figures demonstrate that vanilla-SAC agent performs poorly across nearly the entire defined range, whereas both SAC-R and SAC-BAR outperform significantly. The steady-state errors averaged over the defined range for three algorithms are presented in Table V. Compared to the baseline vanilla-SAC, SAC-R and SAC-BAR achieved reductions in averaged steady-state errors of $27.9\%$ and $36.3\%$, respectively. Specifically, our proposed SAC-BAR reaches the desired orientation at every target angle pair with steady-state errors consistently bounded below $1.5^{\circ}$ except for three outliers. The averaged steady-state error of SAC-BAR is $0.58^{\circ}$, demonstrating its superior control performance in accuracy and robustness across different target angles.
@@
Figure 11: Experimental results in the steady-state error tests of the robotic eye at eyery combination of $(\theta^{*},\psi^{*})$ target angles spaced at $2.5^{\circ}$ intervals for $(|\theta^{*}|,|\psi^{*}|)\leq 10^{\circ}$ where the color bar represents the Euclidean distance norm of steady-state error.
@@
The above experimental results indicate that randomization of muscle dynamics significantly enhances the real-world control performance of the learning-based controllers in the robotic system of interest, while bootstrap and augmentation techniques contribute less to the improvement. We believe that the enhancement observed in SAC-BAR compared to SAC-R primarily stems from the improvement in training efficiency. With the incorporation of bootstrap and augmentation methods, SAC-BAR achieves more comprehensive training and higher averaged rewards in simulation after $3500$ training episodes. This improvement is potentially more significant in scenarios where training data is limited or exploration proves more challenging.
V-C Learning Control of Robotic Eye
@@
To demonstrate the applicability of the proposed learning control framework across diverse robotic systems driven by artificial muscle strings, this subsection presents experimental results on an additional platform, the 2-DOF robotic eye. We focus on the foveation control in the experiment. The robotic eye system, initialized at $\theta=\psi=0^{\circ}$ in ambient room temperature, is configured to foveate at various target angles $(\theta^{*},\psi^{*})$. Considering the system’s rise time, each testing episode lasts for $t_{\text{total}}=15$s, with an action time step of $t_{a}=0.5$s. The averaged Euclidean distance between the current and target angles within the last $5s$ window serves as the steady-state error metric, i.e., $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\psi-\psi^{*})^{2}}$.
@@
Experimental trajectories with targets set at $(\theta^{*},\psi^{*})=(-5^{\circ},7.5^{\circ})$ and $(10^{\circ},-5^{\circ})$ are illustrated in Fig 10. These trajectories demonstrate that the measured angles rise to the target in approximately $5$s, with negligible steady-state error. A visual field test is conducted with a series of target angles spaced at $2.5^{\circ}$ intervals within the defined range $|\theta^{*}|,|\psi^{*}|\leq 10^{\circ}$ to evaluate the overall control performance of the algorithm. The robotic eye is tested to foveate at each target in sequence. The results, illustrated in Fig. 11, reveal excellent control performance, with steady-state errors consistently below $1^{\circ}$ for most target angles.
@@
In prior research [23], we adopted a reinforcement learning-based controller using deep deterministic policy gradient (DDPG) algorithm [50], trained in simulation for $50000$ episodes. Additionally, a reduced-order state observer estimated the real-time temperature of SCP muscles, resulting in satisfactory control performance. In this paper, our proposed algorithm demonstrates comparable control performance with just $2000$ training episodes, representing only $1/25$ of the original requirement. Remarkably, it accomplishes this without the need for real-time muscle temperature information.
VI Conclusions
@@
This paper presented a learning-based control framework designed for the string-type artificial muscle-driven robotic systems, leveraging the SAC algorithm with a GRU hidden layer as the baseline DRL algorithm. To enhance training data efficiency and expedite convergence speed, we proposed two methods including bootstrap by PID and state-based data augmentation. Additionally, we incorporated randomization of artificial muscle dynamic parameters during training to address the sim-to-real gap for enhancing real-world control performance. Extensive experiments and ablation studies on two systems—a 2-DOF robotic eye driven by SCP actuators and a parallel robotic wrist driven by TCA muscles—validated the effectiveness and generalizability of the proposed control framework.
In simulation, the designed bootstrap and augmentation methods significantly reduced the training episodes required for DRL-based controllers. Real-world application experiments demonstrated that randomization of muscle dynamics notably enhanced the control performance of learning-based controllers. Furthermore, our framework exhibited satisfactory performance on both robotic platforms, indicating its great potential of applications in similar systems driven by artificial muscle strings.
Despite the notable performance of the proposed algorithm, several limitations exist. Noticeable jitters in action trajectories may prolong system rise time. Additionally, the configuration of the reward function still has a considerable impact on the experimental outcomes. In future work, we plan to analyze these issues comprehensively using extensive experimental data and refine our algorithms to effectively address these challenges.

### v3

@@ VI-C Ablation Study in Simulation
Training results of our proposed SAC-BAR, with and without bootstrap and augmentation, are presented in Fig. 6(b). As illustrated in the figures, our proposed algorithm exhibits the highest level of data efficiency in both robotic systems. Compared to the SAC algorithm trained solely with dynamics randomization (referred to as “w/o AB”), the integration of bootstrap and augmentation enhances data efficiency and expedites the training process. In the robotic eye system, the baseline algorithm achieves an average reward of $-0.10$ after 2,000 training episodes, whereas our approach achieves comparable performance in only approximately 750 episodes, resulting in a 2.67-fold increase in data efficiency compared to the original level. Similarly, in the robotic wrist system, the baseline algorithm achieves an average reward of $-1.61$ after 3,500 episodes, while ours achieves comparable performance in only about 1,200 episodes, leading to a 2.92-fold increase in data efficiency. Notably, in the case of the more complex robotic wrist system, we find the demonstrations provided by the PID controller yield a more significant acceleration effect during the early stages of training. Meanwhile, the augmentation method demonstrates a favorable impact on enhancing the training speed in both systems and throughout the entire learning process of the algorithms.
VII Real-World Experiments
This section presents the real-world experimental results of the proposed SAC-BAR on two testing robotic platforms. We first demonstrate the results of an ablation study conducted on the robotic wrist platform, which validate the effectiveness of the proposed dynamics randomization in enhancing control performance in real-world scenarios. Next, experimental results on the robotic eye are provided to show the generalizability of SAC-BAR across similar string-type artificial muscle-driven robotic systems. Additionally, we investigate the impact of the randomization range on control performance.
VII-A Ablation Study on Robotic Wrist
@@
Figure 7: Comparison of experimental trajectories between the baseline and SAC-BAR. Target angles are set at ($\theta^{*}=2.5^{\circ},\phi^{*}=-7.5^{\circ}$). (a), (b) The system output angles ($\theta,\phi$) and corresponding actions $a_{t}$ of the baseline SAC algorithm. (c), (d) Those of our proposed SAC-BAR algorithm.
To evaluate the impact of each component on real-world control performances, three different control policies are tested on the robotic wrist platform, including the vanilla SAC algorithm, SAC with only randomization of muscles (SAC-R), and our proposed SAC-BAR. The parallel robotic wrist starts from the initial state with $\theta=\phi=0^{\circ}$ and attempts to reach and maintain the target angle set $(\theta^{*},\phi^{*})$. The action vector $a$ is computed by the trained actors at a time step of $t_{a}=0.5$ s, and each episode spans a duration of $t_{\text{total}}=20$ s. In Fig. 7(a), it is evident that without randomization of muscle dynamics, the vanilla-SAC struggles to achieve the desired orientation. While the agent may eventually reach the desired orientation at certain target angles, maintaining that target angle becomes a challenge, resulting in a considerable steady-state error. Conversely, as observed in Fig. 7(c), the proposed SAC-BAR agent achieves the target state in less than $10$ s and consistently remains within its neighborhood thereafter. As illustrated by the input action trajectories in Fig. 7(d), this success is attributed to the agent learning to maximize the action value during the rising phase and promptly reduce it upon approaching the target angles. Conversely, in Fig. 7(b), there exist serious oscillations in the action trajectories of vanilla-SAC during the initial $10$ s, leading to a prolonged rise time and sometimes even failure to reach the target.
@@
Figure 8: Experimental results of the steady-state error tests in the parallel robotic wrist under selected control algorithms. (a), (b), (c) Heat maps illustrating the experimental results for the SAC, SAC-R, and SAC-BAR, respectively. The combination of target angles $(\theta^{*},\phi^{*})$ is systematically spaced at $2.5^{\circ}$ intervals within the range $(|\theta^{*}|,|\phi^{*}|)\leq 10^{\circ}$. The color bars in the visualization indicate the Euclidean distance norm of steady-state errors. (d) The box charts illustrating the medians, maxima, minima, and quartiles of the steady-state errors for the selected comparative control algorithms.
In order to comprehensively evaluate the control performance of the algorithms across different target angles, we designed a steady-state error field test for the robotic wrist. We selected a series of target angles $(\theta^{*},\phi^{*})$ spaced at $2.5^{\circ}$ intervals within the defined range $\theta^{*},\phi^{*}\in[-10^{\circ},10^{\circ}]$. Each agent was set to run for $25$ s, sequentially targeting each angle pair. The steady-state error is measured and averaged over the last $5$s window, determined by computing the Euclidean distance norm between the current and target angles, defined as $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\phi-\phi^{*})^{2}}$. The steady-state tracking errors for different targets of vanilla-SAC, SAC-R, and SAC-BAR are presented in Fig. 8(a), (b), and (c), respectively. The $\mathrm{x}$ and $\mathrm{y}$ axes therein represent $\theta^{*}$ and $\phi^{*}$, while the color bars represent the averaged steady-state errors. Following that, box charts are illustrated in Fig. 8(d). These figures demonstrate that vanilla-SAC agent underperforms across nearly the entire defined range, whereas both SAC-R and SAC-BAR show significantly improved control performance.
The steady-state errors averaged over the defined range, denoted as $\bar{e}_{\text{ss}}$, for three variants and the PID controller are presented in Table III. In real-world experiments, all three DRL-based controllers demonstrate superior performance compared to the traditional PID controller, reducing the tracking error by $11.6\%$, $35.9\%$, and $43.7\%$, respectively. Compared to the baseline SAC, SAC-R and SAC-BAR achieved reductions in the averaged tracking errors of $27.9\%$ and $36.3\%$, respectively. Specifically, our proposed SAC-BAR reaches the desired orientation at every target angle pair with an averaged error of $0.58^{\circ}$. Furthermore, as shown in Table III, the observed improvement in real-world performance can be attributed to overcoming the sim-to-real gap, as evidenced by the reduced disparity between simulated and actual results.
TABLE III: The experimental (Real) and simulation (Sim) results of steady state error using different control methods
$\bar{e}_{\text{ss}}$ (deg) Control Method
PID SAC SAC-R SAC-BAR
Real $1.03\pm 0.49$ $0.91\pm 0.46$ $0.66\pm 0.39$ $0.58\pm 0.37$
Sim $1.12\pm 0.47$ $0.46\pm 0.27$ $0.47\pm 0.33$ $0.40\pm 0.17$
@@
The above experimental results indicate that the proposed randomization of muscle dynamics significantly enhances the real-world control performance in the robotic systems of interest, while bootstrap and augmentation techniques contribute relatively less to the improvement. We believe that the enhancement observed in SAC-BAR compared to SAC-R primarily stems from the improvement in training efficiency. With the incorporation of bootstrap and augmentation methods, SAC-BAR achieves more comprehensive training and higher averaged rewards in simulation after 3,500 training episodes. We believe this improvement is potentially more significant in scenarios where exploration is more challenging.
VII-B Learning Control of Robotic Eye
Figure 9: Experimental trajectories of the foveation test of the robotic eye. Target angles are set at $(\theta^{*}=-5^{\circ},\psi^{*}=7.5^{\circ})$ and $(\theta^{*}=10^{\circ},\psi^{*}=-5^{\circ})$, respectively. (a), (c) The system output angles $(\theta,\psi)$. (b), (d) The corresponding actions $a_{t}$ generated by the trained actor.
Figure 10: Experimental results on the steady-state tracking errors in robotic eye at eyery combination of $(\theta^{*},\psi^{*})$ target angles spaced at $2.5^{\circ}$ intervals for $(|\theta^{*}|,|\psi^{*}|)\leq 10^{\circ}$ where the color bar represents the Euclidean distance norm of steady-state error.
@@ VII-B Learning Control of Robotic Eye
To demonstrate the applicability of the proposed learning control framework across diverse robotic systems driven by artificial muscle strings, this subsection presents experimental results on an additional platform, the 2-DOF robotic eye. We focus on the foveation control task in the experiment. The robotic eye, initialized at $\theta=\psi=0^{\circ}$ at ambient room temperature, is configured to foveate at various target angles $(\theta^{*},\psi^{*})$. Considering the system’s rise time, each testing episode lasts for $t_{\text{total}}=15$ s, with an action time step of $t_{a}=0.5$ s. The averaged Euclidean distance between the current and target angles within the last $5s$ window serves as the steady-state error metric, i.e., $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\psi-\psi^{*})^{2}}$.
@@
Experimental trajectories with targets set at $(\theta^{*},\psi^{*})=(-5^{\circ},7.5^{\circ})$ and $(10^{\circ},-5^{\circ})$ are illustrated in Fig 9. These trajectories demonstrate that the measured angles rise to the target in approximately $5$ s, with negligible steady-state error. A similar field test is conducted with a series of target angles spaced at $2.5^{\circ}$ intervals within the defined range $|\theta^{*}|,|\psi^{*}|\leq 10^{\circ}$ to evaluate the overall control performance of the algorithm. The robotic eye is tested to foveate at each target in sequence. The results, illustrated in Fig. 10, reveal solid control performance, with steady-state errors consistently below $1^{\circ}$ for most target angles and an average of $0.56^{\circ}$.
@@
In our prior research[20], we employed a DRL-based controller using the Deep Deterministic Policy Gradient algorithm[43], which was trained in simulation for 50,000 episodes. Additionally, a reduced-order state observer was used to estimate the real-time temperature of SCP muscles, achieving satisfactory control performance. In this paper, our proposed algorithm achieves comparable control performance with only 2,000 training episodes, representing just 1/25 of the original training time. Moreover, this is accomplished without the need to estimate real-time muscle temperature information under the condition of ambient temperature remaining relatively stable.
VII-C Randomization Ranges on Control Performance
TABLE IV: Comparison of control errors between models trained under different randomization ranges
$\bar{e}_{\text{ss}}$ (deg) Randomization Range
Small Original Large
Real $0.63\pm 0.49$ $0.58\pm 0.37$ $0.76\pm 0.43$
Sim $0.38\pm 0.17$ $0.40\pm 0.17$ $0.74\pm 0.41$
In designing dynamics randomization, maintaining a balance between randomization level and model fidelity is crucial. Excessive randomization leads to an overly conservative controller, which may over-generalize and fail to leverage the more predictable aspects of system dynamics. Conversely, insufficient randomization causes the controller to overfit to a narrow set of conditions, thus limiting its performance when deployed in real-world systems. To investigate this trade-off, we conducted an experiment in which our control framework was trained with two additional randomization ranges—“small” and “large.” The small range was defined as half of the original randomization range, while the large range was doubled. We evaluated the performance of controllers trained under these varying randomization ranges in both simulated and real-world environments on the robotic wrist platform. The results, as summarized in Table IV, indicate that the control performance experiences a significant decline when randomization range is excessively large. Conversely, the smaller randomization range impairs the controller’s generalization capability from simulation to reality. Overall, a balanced range proves essential for ensuring robust performance.
VIII Conclusion
@@
This paper presents a learning-based control framework SAC-BAR designed for the string-type artificial muscle-driven robotic systems. To enhance training data efficiency and expedite convergence speed, we proposed two methods, namely bootstrap via PID and state-based data augmentation. Additionally, we incorporated randomization of artificial muscle dynamic parameters during training to address the sim-to-real gap for enhancing real-world control performance. Extensive experiments and ablation studies on two systems—a robotic eye driven and a parallel robotic wrist—validated the effectiveness and generalizability of the proposed control framework. In simulation, the designed bootstrap and augmentation methods reduced the required training episodes by approximately 60% compared to the baseline algorithm. Real-world experiments demonstrated that randomization of muscle dynamics improved the control performance of DRL-based controllers in targeted robotic systems, reducing the average error by 36.3%. Furthermore, our framework showed satisfactory performance on both robotic platforms, highlighting its potential applicability to similar artificial muscle robots.
Despite the strong performance of the proposed SAC-BAR, several limitations exist. The configuration of the reward function has a significant impact on experimental outcomes, often requiring empirical tuning to ensure optimal controller performance. The dynamic models of the artificial muscles and robotic wrist are simplified and do not fully represent their true dynamics. In future work, we plan to explore automated methods for reward function tuning, potentially incorporating adaptive or self-tuning mechanisms that respond dynamically to performance metrics. To better account for real-world variability, we will extend the randomization of muscle dynamics by incorporating high-fidelity models that capture nonlinearity and time-dependent factors, using data-driven or hybrid modeling approaches.